The cost of attending hope university for one year is $60000 for tuition, $20000 for room accomodatio,$15000 for meal and $5000 for book andbsupplies. The student could also have earned $150000 by getting a job instead of going to a university and 10% interest on expenses he or she incured at yhe begining of the year. You are required to calculate the.Explicit cost,implicit cost and total economy cost of attending the unversity.
1
Expert's answer
2019-09-02T09:07:20-0400

Explicit cost = 60,000 + 20,000 + 15,000 + 5,000 = 100,000.

Implicit cost = 150,000 + 100,000×0.1 = 160,000.

Total economy cost of attending the unversity = Explicit cost + Implicit cost = 100,000 + 160,000 = 260,000.